# Learning Unknown Intervention Targets in Structural Causal Models from Heterogeneous Data

### 11 Dec 2023 v1

@@ 1 Introduction
Causal relationships among a set of variables in a system can be modeled by a structural causal model (SCM) where each variable is a function of its direct causes and some exogenous noise. An intervention on a variable can be considered as modifying its causal mechanism, i.e., changing the conditional probability distribution of the intervened variable given its direct causes. In randomized control trials, randomized interventions on a target variable are utilized to estimate the causal effect of the target. However, in some applications, we may not have full control in terms of which variables are intervened on. For instance, in recovering causal protein-signaling networks from single-cell data (Sachs et al.,, 2005; Ness et al.,, 2017), drugs are injected into cells to inhibit or activate some signaling proteins, and gene expression levels are measured. In these experiments, the intervention targets are unknown. Moreover, in some cases, an intervention is done by an unknown source and we must locate the source of the intervention in the system. As an example, microservices systems in cloud clusters are vulnerable to faults such as equipment failures or adversarial attacks. It is crucial to locate the root cause of faulty operation in the system by identifying the source of fault/intervention (Aggarwal et al.,, 2021; Budhathoki et al.,, 2022). In these examples, the collected data is often heterogeneous and is gathered from multiple domains/environments where the causal mechanisms of some of the variables are changing across the environments.
@@
    For the recovery phase, we provide identifiability results for recovering the exogenous noises whose distributions change across the environments. In particular, in nonlinear causal models with exogenous noises belonging to an exponential family, the recovery is possible under some mild invertibility assumption (Assumption 1(a)) for causally sufficient systems (i.e., there are no latent confounders). For systems with latent variables, under some further assumptions (Assumption 1(b)), we show that the recovery is still possible (Proposition 1).
@@
    For the matching phase, we prove that LIT algorithm recovers the true intervention targets for causally sufficient systems using the recovered exogenous noises (Theorem 1). LIT algorithm requires only quadratic number of CI tests while previous work (Jaber et al.,, 2020) requires exponential number of CI/invariance tests with respect to number of variables in the system. In the presence of latent confounders, we show that LIT algorithm returns a superset of true intervention targets and present a graphical characterization for the recovery output (Theorem 2). Unlike previous work, LIT algorithm allows for latent confounders to change across environments. Moreover, for the setting studied in the literature (i.e., all latent confounders are not changing across environments), our recovery output is more informative than the state-of-the-art (see Remark 2 for more details).
@@
    Experimental results show that LIT outperforms previous work in recovering intervention targets in the presence of latent confounders or when the underlying SCM is nonlinear.
@@ 3.2 Recovery Phase
As we have access to a collection of distributions in $\mathcal{P}$, we will exploit the heterogeneity in data to recover the exogenous noises in ${\bf N}_{{\bf T}}$. Specifically, we utilize a contrastive learning approach. We observe auxiliary variable $U$ indicating the index of the environment, and train a nonlinear regression model with universal approximation capability [^3] Universal approximation capability refers to the ability to approximate any Borel measurable function to any desired degree of accuracy. See (Hornik et al.,, 1989) for more detail. for the supervised learning task (Please refer to Appendix A.1 for more details). We consider a specific exponential family for the distributions of the exogenous noises (see Appendix A.1 for the definition of the exponential family and the assumptions on it). To ensure that the noises in ${\bf N}_{\bf T}$ can be recovered, we require the following assumption:
@@
Assumption 1(a) is standard in contrastive learning algorithms under causal sufficiency assumption. It is satisfied for all acyclic linear SCMs and nonlinear additive noise models. To extend the results to the latent confounder setting, we added Assumption 1(b). In particular, vector ${\bf V}$ is the recovered part that is invariant across the environments and is independent of ${\bf N}_{{\bf T}}$ given the index of the environment. It corresponds to a function of exogenous noises whose distributions do not change across the environments.
@@
Please refer to Appendix A for a detailed description of the contrastive learning approach and extra discussion about when Assumption 1 is satisfied.
@@ 3.3 Matching phase
Throughout the matching phase, we assume that the exogenous noises in ${\bf N}_{{\bf T}}$ are recovered up to some permutation and component-wise invertible transformations. Denote the recovered noise corresponding to the noise $N_{i}\in{\bf N}_{\bf T}$ as $\tilde{N}_{i}$ (i.e., $\tilde{N}_{i}$ is an invertible transformation of $N_{i}$), and denote the collection of all the recovered noises as $\tilde{{\bf N}}_{\bf T}$. Note that we cannot learn the correspondence of the recovered noises to the true noises due to permutation indeterminacy according to Proposition 1. In fact, the goal of the matching phase is to recover the mapping between the recovered noises in $\tilde{{\bf N}}_{\bf T}$ and the corresponding exogenous noises in ${\bf N}_{\bf T}$. [^4] Due to the permutation indeterminacy, there exists a one-to-one mapping $\sigma$ that maps each noise in ${\bf N}_{\bf T}$ to a distinct noise in $\tilde{{\bf N}}_{\bf T}$. For notation simplicity, we denote $\sigma(N_{i})\in\tilde{{\bf N}}_{\bf T}$ as $\tilde{N}_{i}$ for each $N_{i}\in{\bf N}_{\bf T}$. Since each noise $N_{i}\in{\bf N}_{\bf T}$ corresponds to a variable $X_{i}\in{\bf T}$, the goal of the matching phase is to learn the inverse mapping $\sigma^{-1}$ that maps the recovered noises to the variables in ${\bf T}$.
@@ 3.3.1 T-faithfulness assumption
The model is T-faithful to the augmented graph $G_{{\bf T}}$, in the sense that for any noise $N_{i}\in{\bf N}_{\bf T}$, observed variable $X_{k}\in{\bf O}$, and disjoint sets ${\bf W}_{1}\subseteq{\bf O}\setminus\{X_{k}\}$, ${\bf W}_{2}\subseteq{\bf N}_{\bf T}\setminus\{N_{i}\}$: $(N_{i}\perp\!\!\!\perp X_{k}|{\bf W}_{1},{\bf W}_{2})_{G_{\bf T}}$ if and only if $\tilde{N}_{i}\perp\!\!\!\perp X_{k}|{\bf W}_{1},\tilde{{\bf W}}_{2}$, where $\tilde{N}_{i}$ is an arbitrary invertible transformation of $N_{i}$, and $\tilde{{\bf W}}_{2}=\{\tilde{N}_{j}|N_{j}\in{\bf W}_{2}\}$.
@@ 3.3.2 Matching phase under causal sufficiency
For each variable $X_{i}$, define the possible parent set ${\bf S}_{i}$ as the set of variables whose indicator set is a strict subset of ${\bf I}_{i}$, i.e., ${\bf S}_{i}:=\{X_{j}|{\bf I}_{j}\subsetneq{\bf I}_{i},1\leq j\leq n\}$. Define the residual set $\mathcal{N}_{i}$ as the set of noises in ${\bf I}_{i}$ that do not belong to any indicator set of the variables in ${\bf S}_{i}$, i.e. $\mathbf{I}_{i}\setminus\cup_{j:X_{j}\in{\bf S}_{i}}{\bf I}_{j}$. ${\bf S}_{i}$ includes a subset of the ancestors of $X_{i}$, and no descendants of $X_{i}$ are included in ${\bf S}_{i}$. Further, $\mathcal{N}_{i}$ represent the noises in $X_{i}$ that do not affect $X_{i}$ through variables in ${\bf S}_{i}$. Under causal sufficiency assumption, $|\mathcal{N}_{i}|$ is either 0 or 1 (see Appendix D.3). The following proposition provides the conditions for checking whether a variable $X_{i}$ is in the intervention target set or not given the indicator sets $\mathcal{I}$ and ${\bf S}_{i}$. Equipped with this proposition, we devise LIT Algorithm (see Algorithm 1) which recovers ${\bf T}$ under the causal sufficiency assumption.
@@
    If $\mathcal{N}_{i}\neq\emptyset$ and ${\bf I}_{i}$ is not unique, let $\mathcal{X}_{i}=\{X_{i_{1}},\cdots,X_{i_{p}}\}$, $p\geq 2$, be the set of all variables with the same indicator set as $X_{i}$ including $X_{i}$ itself. [^5] As all the variables in $\mathcal{X}_{i}$ have the same indicator set, their corresponding possible parent sets ${\bf S}_{i_{k}}$ and residual sets $\mathcal{N}_{i_{k}}$ are also equal. In the statement of the proposition, we use ${\bf I}_{i}$, ${\bf S}_{i}$, $\mathcal{N}_{i}$ to denote the indicator set, possible parent set and residual set corresponding to any variable in $\{X_{i_{1}},\cdots,X_{i_{p}}\}$. Suppose $\mathcal{N}_{i}=\{\tilde{N}_{l}\}$ for some $\tilde{N}_{l}\in\tilde{{\bf N}}_{\bf T}$. Then, the variable $X_{i_{k}}$ satisfying the following condition is the only variable from $\mathcal{X}_{i}$ that is in ${\bf T}$ , i.e., $X_{i_{k}}\in{\bf T}$ if all other variables in $\mathcal{X}_{i}$ are independent of $\tilde{N}_{l}$ conditioned on $X_{i_{k}}$ and ${\bf S}_{i}$.
@@
Recall that one observed variable is in the intervention target set ${\bf T}$ if and only if its corresponding exogenous noise is recovered in $\tilde{{\bf N}}_{\bf T}$. The statement in I holds because if $X_{i}\in{\bf T}$, then $\tilde{N}_{i}$ cannot appear in ${\bf I}_{j}$ for any non-descendant $X_{j}$ of $X_{i}$. When $\mathcal{N}_{i}\neq\emptyset$, this means that the only noise $\tilde{N}_{l}\in\mathcal{N}_{i}$ is either the exogenous noise of $X_{i}$, or the exogenous noise of some ancestor of $X_{i}$ whose indicator set is the same as ${\bf I}_{i}$. We can then use conditions II and 5 to further distinguish between these two cases. If there are no other variables with the same indicator set (i.e., ${\bf I}_{i}$ is unique), then $X_{i}\in{\bf T}$. Otherwise, among the variables with the same indicator set, there is only one variable $X_{i_{k}}$ that corresponds to $\tilde{N}_{l}$ and belongs to the intervention target set, and the rest are descendants of $X_{i_{k}}$. Herein, we use (C1) to find such $X_{i_{k}}$, as given $X_{i_{k}}$ and ${\bf S}_{i}$, $\tilde{N}_{l}$ becomes independent from $X_{i_{j}}$ for all $j\neq k$ under T-faithfulness assumption. Please refer to Appendix B.2 for an example explaining how the aforementioned three conditions can be used to recover ${\bf T}$.
Based on Proposition 2, we propose LIT algorithm (see Algorithm 1) which returns a candidate intervention set ${\bf K}$. Specifically, we first check if a variable can be added to or excluded from ${\bf K}$ according to I and II. We then partition the remaining variables in ${\bf U}$ into disjoint subsets (which correspond to the collection of all $\mathcal{X}_{i}$ in 5), and find the candidate in each set (denoted by ${\bf K}_{{\bf U}_{i}}$) using condition 5. Note that LIT algorithm only requires quadratic number of CI tests: $O(n|{\bf T}|)$ for constructing the indicator set, and at most $O(n|{\bf T}|)$ for checking (C1). This is a significant reduction from the exponential number of independence/invariance tests with respect to $n$ in the literature (Jaber et al.,, 2020; Mooij et al.,, 2020).
@@ 3.3.3 Matching phase in the presence of latent confounders
    $X_{i}\not\in{\bf T_{O}}$ if $\mathcal{N}_{i}\neq\emptyset$, and every recovered noise in the residual set $\mathcal{N}_{i}$ belongs to at least one other indicator set ${\bf I}_{j_{l}}$, where ${\bf I}_{j_{l}}$ is not a strict superset of ${\bf I}_{i}$:
@@
[…]
    $\tilde{N}_{l}\perp\!\!\!\perp X_{i_{j}}|\left(\cup_{k^{\prime}\in K}X_{i_{k^{\prime}}}\right)\cup{\bf S}\cup\left({\bf I}_{i}\setminus\{\tilde{N}_{l}\}\right).$
The statement in IV holds because if $X_{i}\in{\bf T_{O}}$, then its corresponding exogenous noise $\tilde{N}_{i}$ must be in $\mathcal{N}_{i}$. Any variable (such as $X_{j_{l}}$) that is dependent on $\tilde{N}_{l}$ must be a descendant of $X_{i}$, and hence have ${\bf I}_{i}\subseteq{\bf I}_{j_{l}}$. For C2, similar to the argument for condition 5, there is at most one variable (say $X_{i_{k}}$) in $\mathcal{X}_{i}$ that belongs to ${\bf T_{O}}$. Moreover, if $X_{i_{k}}\in{\bf T_{O}}$, then all other variables in $\mathcal{X}_{i}$ are its descendants. The recovered noises in $\mathcal{N}_{i}$ cannot be conditionally independent of $X_{i_{k}}$, as they correspond to either $X_{i_{k}}$ or some latent confounder that is a parent of $X_{i_{k}}$. Therefore, if an observed variable $X_{i_{j}}\in\mathcal{X}_{i}$ is conditionally independent of a recovered noise in $\mathcal{N}_{i}$ given some other variables in the system, then it cannot be an intervention target. Note that under the causal sufficiency assumption, $X_{i_{k}}$ and ${\bf S}_{i}$ are sufficient for the conditioning set. Therefore condition C2 reduces to 5. When latent confounders are present, $X_{i_{k}}$ and ${\bf S}_{i}$ may not be sufficient. However, condition C2 states that in order to perform such a CI test, it suffices to consider subsets of $\mathcal{X}_{i}\backslash\{X_{i_{j}}\}$ and ${\bf S}_{i}$ in the conditioning set as their union contains all the ancestors of $X_{i_{j}}$ among the observed variables.
Based on Proposition 3, we update the LIT algorithm in the presence of latent confounders. We check condition IV in line 4, and replace condition 5 by C2 in line 8. We keep line 5 in the latent case. In fact, if $X_{i}$ is not ruled out by conditions I and IV, it is added to ${\bf K}$ if ${\bf I}_{i}$ is unique as it cannot be ruled out by condition C2 either. However, the uniqueness of the indicator set does not necessarily imply that the variable belongs to ${\bf T_{O}}$. Lastly, note that under causal sufficiency assumption, condition IV is automatically satisfied, and condition C2 reduces to condition 5. Hence the algorithm remains consistent with the causally sufficient case.
@@
Given a SCM and its corresponding augmented graph $G_{\bf T}$, the auxiliary graph $Aux(G_{\bf T})$ is constructed from $G_{\bf T}$ as follows:
@@
    For each $X_{i}\in{\bf T_{O}}$ with its corresponding exogenous noise $N_{i}$, add the edge $N_{i}\rightarrow X_{j}$ if (i) there is an edge from $N_{i}$ to $X_{j}$ in the MAG corresponding to $G_{\bf T}$, and (ii) $An_{G_{\bf T}}(X_{i})\cap{\bf T}=An_{G_{\bf T}}(X_{j})\cap{\bf T}$ [^6] Recall from Section 3.3.1 that under Assumption 2, $An_{G_{\bf T}}(X_{i})\cap{\bf N}_{\bf T}$ can be recovered from ${\bf I}_{i}$..
@@
    For each $N_{l}\in{\bf N}_{{\bf T_{L}}}$ (noise that corresponds to a variable in ${\bf T_{L}}$), add or remove the edge from $N_{l}$ to observed variables $X_{k}$ according to some extra conditions (see Appendix B.3 for formal statement).
@@
We can make the following observations from Theorem 2. First, under the causal sufficiency assumption, Theorem 2 implies that ${\bf T_{O}}$ can be uniquely identified. This is because no edges are added to $Aux(G_{\bf T})$ compared with $G_{\bf T}$ according to Definition 1. Second, if all latent variables are not intervention targets (i.e., ${\bf T}={\bf T_{O}}$), then our identifiability result is stronger than the existing results in (Jaber et al.,, 2020; Varici et al.,, 2022). In particular, they showed that in this case, ${\bf T_{O}}$ can only be identified up to the neighbors of ${\bf N}_{{\bf T}}$ in the MAG corresponding to $G_{\bf T}$. We improve their results by adding part (a)(ii) in Definition 1, due to the recovery of $\tilde{{\bf N}}_{{\bf T}}$. See Example 2 below.

### 09 Mar 2024 v2

@@ 1 Introduction
Causal relationships among a set of variables in a system can be modeled by a structural causal model (SCM) where each variable is a function of its direct causes and some exogenous noise. An intervention on a variable can be considered as modifying its causal mechanism, i.e., changing the conditional probability distribution of the intervened variable given its direct causes. In randomized control trials, randomized interventions on a target variable are utilized to estimate the causal effect of the target. However, in some applications, we may not have full control in terms of which variables are intervened on. For instance, in recovering causal protein-signaling networks from single-cell data (Sachs et al.,, 2005; Ness et al.,, 2017), drugs are injected into cells to inhibit or activate some signaling proteins, and gene expression levels are measured. In these experiments, the intervention targets are unknown. Moreover, in some cases, an intervention is done by an unknown source and we must locate the source of the intervention in the system. As an example, microservices systems in cloud clusters are vulnerable to faults such as equipment failures or adversarial attacks. It is crucial to locate the root cause of faulty operation in the system by identifying the source of fault/intervention (Aggarwal et al.,, 2021; Budhathoki et al.,, 2022; Ikram et al.,, 2022). In these examples, the collected data is often heterogeneous and is gathered from multiple domains/environments where the causal mechanisms of some of the variables are changing across the environments.
@@
    For the recovery phase, we provide identifiability results for recovering the exogenous noises whose distributions change across the environments. In particular, in nonlinear causal models with exogenous noises from an exponential family, the recovery is possible under some mild invertibility assumption (Assumption 1(a)) for causally sufficient systems (i.e., there are no latent confounders). For systems with latent confounders, under some further assumptions (Assumption 1(b)), we show that the recovery is still possible (Proposition 1).
@@
    For the matching phase, we prove that LIT algorithm recovers the true intervention targets for causally sufficient systems using the recovered exogenous noises (Theorem 1). LIT algorithm requires merely quadratic number of CI tests. In comparison, previous work of (Jaber et al.,, 2020) required an exponential number of CI/invariance tests with respect to number of variables in the system. In the presence of latent confounders, we show that LIT algorithm returns a superset of true intervention targets and present a graphical characterization of the output of recovery (Theorem 2). Unlike previous work, LIT algorithm allows latent confounders to change across environments. Moreover, for the setting studied in the literature (i.e., when latent confounders do not change across environments), our recovery output is more informative than the state-of-the-art (see Remark 3 for more details).
@@
    Our experimental results show that LIT outperforms previous work in recovering intervention targets in the presence of latent confounders as well as when the underlying SCM is nonlinear.
@@ 3.2 Recovery Phase
As we have access to a collection of distributions in $\mathcal{P}$, we will exploit the heterogeneity in data to recover the exogenous noises in ${\bf N}_{{\bf T}}$. Specifically, we utilize a contrastive learning approach. We observe auxiliary variable $U$ indicating the index of the environment, and train a nonlinear regression model with universal approximation capability [^3] Universal approximation capability refers to the ability to approximate any Borel measurable function to any desired degree of accuracy. See (Hornik et al.,, 1989) for more detail. for the supervised learning task (Please refer to Appendix A for more details). We consider an exponential family for the distributions of the exogenous noises (see Appendix A for the definition). To ensure that the noises in ${\bf N}_{\bf T}$ can be recovered, we require the following assumption:
@@
Assumption 1(a) is standard in contrastive learning algorithms under the causal sufficiency assumption. It is satisfied for all acyclic linear SCMs and nonlinear additive noise models. To extend the results to the latent confounder setting, we added Assumption 1(b). In particular, vector ${\bf V}$ is the recovered part that is invariant across the environments and is independent of ${\bf N}_{{\bf T}}$ given the index of the environment. It corresponds to a function of exogenous noises whose distributions do not change across the environments.
@@
Please refer to Appendix A for a detailed description of the contrastive learning approach and extra discussions about when Assumption 1 is satisfied.
Remark 1.
Note that the assumption on the number of environments $D$ is required for recovering the noises in $\mathbf{N}_{\mathbf{T}}$ without any prior knowledge of $f_{i}$’s functional form. We can relax the assumptions on the number of environments by adding restrictions on $f_{i}$ (or equivalently on $\mathbf{g}_{\mathcal{M}}$). For example, for piecewise linear $\mathbf{g}_{\mathcal{M}}$, under additional assumptions on the exogenous noises, we can recover $\mathbf{N}_{\mathbf{T}}$ from merely two environments in causally sufficient cases, based on the results in (Kivva et al.,, 2022).
@@ 3.3 Matching Phase
Throughout the matching phase, we assume that the exogenous noises in ${\bf N}_{{\bf T}}$ are recovered up to some permutation and component-wise invertible transformations. Denote the recovered noise corresponding to the noise $N_{i}\in{\bf N}_{\bf T}$ as $\tilde{N}_{i}$ (i.e., $\tilde{N}_{i}$ is an invertible transformation of $N_{i}$), and denote the collection of all the recovered noises as $\tilde{{\bf N}}_{\bf T}$. Note that we cannot learn the correspondence of the recovered noises to the true noises due to permutation indeterminacy according to Proposition 1. However, since a variable $X_{i}$ is in ${\bf T}$ if and only if its exogenous noise $N_{i}\in{\bf N}_{\bf T}$, we can recover the intervention targets by matching the observed variables to the recovered noises. [^4] Due to the permutation indeterminacy, there exists a one-to-one mapping $\sigma$ that maps each noise in ${\bf N}_{\bf T}$ to a distinct noise in $\tilde{{\bf N}}_{\bf T}$. For notation simplicity, we denote $\sigma(N_{i})\in\tilde{{\bf N}}_{\bf T}$ as $\tilde{N}_{i}$ for each $N_{i}\in{\bf N}_{\bf T}$. The matching phase aims to match each recovered noise to its corresponding observed variable (if such a variable exists). Note that a recovered noise may correspond to a latent confounder. Further, an observed variable does not correspond to any recovered noise if it is not in the intervention target set.
@@ 3.3.1 T-faithfulness assumption
The model is T-faithful to the augmented graph $G_{{\bf T}}$, in the sense that for any noise $N_{i}\in{\bf N}_{\bf T}$, observed variable $X_{k}\in{\bf O}$, and disjoint sets ${\bf W}_{1}\subseteq{\bf O}\setminus\{X_{k}\}$, ${\bf W}_{2}\subseteq{\bf N}_{\bf T}\setminus\{N_{i}\}$: $(N_{i}\perp\!\!\!\perp X_{k}|{\bf W}_{1},{\bf W}_{2})_{G_{\bf T}}$ if and only if $\tilde{N}_{i}\perp\!\!\!\perp X_{k}|{\bf W}_{1},\tilde{{\bf W}}_{2}$, where $\tilde{N}_{l}$ is an arbitrary invertible transformation of $N_{l}$ for any $N_{l}\in{\bf N}_{\bf T}$, and $\tilde{{\bf W}}_{2}=\{\tilde{N}_{j}|N_{j}\in{\bf W}_{2}\}$.
@@ 3.3.2 Matching phase under causal sufficiency
For each variable $X_{i}$, define the possible parent set ${\bf S}_{i}$ as the set of variables whose indicator set is a strict subset of ${\bf I}_{i}$, i.e., ${\bf S}_{i}:=\{X_{j}|{\bf I}_{j}\subsetneq{\bf I}_{i},1\leq j\leq n\}$. Define the residual set $\mathcal{N}_{i}$ as the set of noises in ${\bf I}_{i}$ that do not belong to any indicator set of the variables in ${\bf S}_{i}$, i.e. $\mathbf{I}_{i}\setminus\cup_{j:X_{j}\in{\bf S}_{i}}{\bf I}_{j}$. ${\bf S}_{i}$ includes a subset of the ancestors of $X_{i}$, and no descendants of $X_{i}$ are included in ${\bf S}_{i}$. Further, $\mathcal{N}_{i}$ represent the recovered noises in $X_{i}$ that do not affect $X_{i}$ through variables in ${\bf S}_{i}$. Under the causal sufficiency assumption, $|\mathcal{N}_{i}|$ is either 0 or 1 (see Appendix D.3). The following proposition provides the conditions for checking whether a variable $X_{i}$ is in the intervention target set or not given the indicator sets $\mathcal{I}$ and ${\bf S}_{i}$. Equipped with this proposition, we devise LIT Algorithm (Algorithm 1) which recovers ${\bf T}$ under the causal sufficiency assumption.
@@
    If $\mathcal{N}_{i}\neq\emptyset$ and ${\bf I}_{i}$ is not unique, let $\mathcal{X}_{i}=\{X_{i_{1}},\cdots,X_{i_{p}}\}$, $p\geq 2$, be the set of all variables with the same indicator set as $X_{i}$, including $X_{i}$ itself. [^5] As all the variables in $\mathcal{X}_{i}$ have the same indicator set, their corresponding possible parent sets ${\bf S}_{i_{k}}$ and residual sets $\mathcal{N}_{i_{k}}$ are also equal. In the statement of the proposition, we use ${\bf I}_{i}$, ${\bf S}_{i}$, $\mathcal{N}_{i}$ to denote the indicator set, possible parent set and residual set corresponding to any variable in $\{X_{i_{1}},\cdots,X_{i_{p}}\}$. Suppose $\mathcal{N}_{i}=\{\tilde{N}_{l}\}$ for some $\tilde{N}_{l}\in\tilde{{\bf N}}_{\bf T}$. Then, the variable $X_{i_{k}}$ satisfying the following condition is the only variable from $\mathcal{X}_{i}$ that is in ${\bf T}$ , i.e., $X_{i_{k}}\in{\bf T}$ if and only if all other variables in $\mathcal{X}_{i}$ are independent of $\tilde{N}_{l}$ conditioned on $X_{i_{k}}$ and ${\bf S}_{i}$.
@@
Recall that an observed variable is in the intervention target set ${\bf T}$ if and only if its corresponding exogenous noise is recovered in $\tilde{{\bf N}}_{\bf T}$. The statement in I holds because if $X_{i}\in{\bf T}$, then $\tilde{N}_{i}$ cannot appear in ${\bf I}_{j}$ for any non-descendant $X_{j}$ of $X_{i}$. When $\mathcal{N}_{i}\neq\emptyset$, this means that the only noise $\tilde{N}_{l}\in\mathcal{N}_{i}$ is either the exogenous noise of $X_{i}$, or the exogenous noise of some ancestor of $X_{i}$ whose indicator set is the same as ${\bf I}_{i}$. We can then use conditions II and 5 to further distinguish between these two cases. If there are no other variables with the same indicator set (i.e., ${\bf I}_{i}$ is unique), then $X_{i}\in{\bf T}$. Otherwise, among the variables with the same indicator set, there is only one variable $X_{i_{k}}$ that corresponds to $\tilde{N}_{l}$ and belongs to the intervention target set, and the rest are descendants of $X_{i_{k}}$. Herein, we use (C1) to find such $X_{i_{k}}$, as given $X_{i_{k}}$ and ${\bf S}_{i}$, $\tilde{N}_{l}$ becomes independent from $X_{i_{j}}$ for all $j\neq k$ under T-faithfulness assumption. Please refer to Appendix B.2 for an example explaining how the aforementioned three conditions can be used to recover ${\bf T}$.
@@
Based on Proposition 2, we propose LIT algorithm (see Algorithm 1) which returns a candidate intervention set ${\bf K}$. Specifically, we first check if a variable can be added to or excluded from ${\bf K}$ according to I and II. We then partition the remaining variables in ${\bf U}$ into disjoint subsets (which correspond to the collection of all $\mathcal{X}_{i}$ in 5), and find the candidate in each set (denoted by ${\bf K}_{{\bf U}_{i}}$) using condition 5. Note that LIT algorithm only requires quadratic number of CI tests: $O(n|{\bf T}|)$ for constructing the indicator set, and at most $O(n^{2}|{\bf T}|)$ for checking (C1). This is a significant reduction from the exponential number of independence/invariance tests with respect to $n$ in the literature (Jaber et al.,, 2020; Mooij et al.,, 2020).
@@ 3.3.3 Matching phase in the presence of latent confounders
    $X_{i}\not\in{\bf T_{O}}$ if $\mathcal{N}_{i}\neq\emptyset$ and every recovered noise in the residual set $\mathcal{N}_{i}$ belongs to at least one other indicator set ${\bf I}_{j_{l}}$ for some $X_{j_{l}}\in{\bf O}$, where ${\bf I}_{j_{l}}$ is not a superset of ${\bf I}_{i}$:
@@
    $\exists K\subseteq[p]\setminus\{j\},~{}\exists{\bf S}\subseteq{\bf S}_{i},~{}\exists\tilde{N}_{l}\in\mathcal{N}_{i}\text{ s.t. }\tilde{N}_{l}\perp\!\!\!\perp X_{i_{j}}|\left(\cup_{k^{\prime}\in K}X_{i_{k^{\prime}}}\right)\cup{\bf S}\cup\left({\bf I}_{i}\setminus\{\tilde{N}_{l}\}\right).$ (C2)
The statement in IV holds because if $X_{i}\in{\bf T_{O}}$, then its corresponding exogenous noise $\tilde{N}_{i}$ must be in $\mathcal{N}_{i}$. Any variable (such as $X_{j_{l}}$) that is dependent on $\tilde{N}_{i}$ must be a descendant of $X_{i}$, and hence have ${\bf I}_{i}\subseteq{\bf I}_{j_{l}}$. For C2, similar to the argument for condition 5, there is at most one variable (say $X_{i_{k}}$) in $\mathcal{X}_{i}$ that belongs to ${\bf T_{O}}$. Moreover, if $X_{i_{k}}\in{\bf T_{O}}$, then all other variables in $\mathcal{X}_{i}$ are its descendants. The recovered noises in $\mathcal{N}_{i}$ cannot be conditionally independent of $X_{i_{k}}$, as they correspond to either $X_{i_{k}}$ or some latent confounder that is a parent of $X_{i_{k}}$. Therefore, if an observed variable $X_{i_{j}}\in\mathcal{X}_{i}$ is conditionally independent of a recovered noise in $\mathcal{N}_{i}$ given some other variables in the system, then it cannot be an intervention target. Note that under the causal sufficiency assumption, $X_{i_{k}}$ and ${\bf S}_{i}$ are sufficient for the conditioning set. Therefore condition C2 reduces to 5. When latent confounders are present, $X_{i_{k}}$ and ${\bf S}_{i}$ may not be sufficient. However, condition C2 states that in order to perform such a CI test, it suffices to consider subsets of $\mathcal{X}_{i}\backslash\{X_{i_{j}}\}$ and ${\bf S}_{i}$ in the conditioning set as their union contains all the ancestors of $X_{i_{j}}$ among the observed variables.
Based on Proposition 3, we update the LIT algorithm in the presence of latent confounders. We check condition IV in line 4, and replace condition 5 by C2 in line 8. We keep line 5 in the latent case. In fact, if $X_{i}$ is not ruled out by conditions I and IV, it is added to ${\bf K}$ if ${\bf I}_{i}$ is unique as it cannot be ruled out by condition C2 either. However, the uniqueness of the indicator set does not necessarily imply that the variable belongs to ${\bf T_{O}}$. Lastly, note that under the causal sufficiency assumption, condition IV is automatically satisfied, and condition C2 reduces to condition 5. Hence the algorithm remains consistent with the causally sufficient case.
@@
For each observed variable $X_{i}$, denote ${\bf I}_{0}(X_{i})$ as $An_{G_{\bf T}}(X_{i})\cap{\bf N}_{\bf T}$. Given an SCM and its corresponding augmented graph $G_{\bf T}$, the auxiliary graph $Aux(G_{\bf T})$ is constructed from $G_{\bf T}$ as follows:
@@
    (i) For each $\tilde{N}_{l}\in\tilde{{\bf N}}_{{\bf T_{L}}}$ (noise that corresponds to a variable in ${\bf T_{L}}$) and each of its child $X_{i}$, keep the edge $\tilde{N}_{l}\rightarrow X_{i}$ if for any other child $X_{j}$ of $\tilde{N}_{l}$ in $G_{\bf T}$, ${\bf I}_{0}(X_{i})\subseteq{\bf I}_{0}(X_{j})$. Otherwise remove the edge $\tilde{N}_{l}\rightarrow X_{i}$. (ii) For any remaining edge $\tilde{N}_{l}\rightarrow X_{i}$, add (remove) the edge $\tilde{N}_{l}\rightarrow X_{k}$ if there is an (no) inducing path between $X_{k}$ and all (some) parents of $X_{i}$ in $\tilde{{\bf N}}_{{\bf T_{L}}}$ relative to ${\bf L}\setminus{\bf T_{L}}$ in $G_{\bf T}$, and ${\bf I}_{0}(X_{k})={\bf I}_{0}(X_{i})$.
Recall from Section 3.3.1 that under Assumption 2, $An_{G_{\bf T}}(X_{i})\cap{\bf N}_{\bf T}$ can be recovered by ${\bf I}_{i}$. Therefore we use ${\bf I}_{0}(X_{i})$ to represent the true value of ${\bf I}_{i}$ which does not depend on the recovery output. Further, it is noteworthy that while part (b)(ii) in Definition 1 includes both adding and removing of the edges, the addition or removal of edges in the step is independent of the order of the edge selection $\tilde{N}_{l}\rightarrow X_{i}$.
@@
Previous work (Jaber et al., (2020)) LIT Algorithm (Theorem 2) True interven- tion target
Figure 2: Comparison of the recovery outputs when latent confounders are not in ${\bf T}$. The dashed circle represents the theoretical limitation of the recovery of ${\bf T}$. It may include observed variables that are theoretically indistinguishable from the true intervention targets due to the presence of latent confounders.
@@
We can draw the following observations from Theorem 2. First, under the causal sufficiency assumption, Theorem 2 implies that ${\bf T_{O}}$ can be uniquely identified. This is because no edges are added to $Aux(G_{\bf T})$ compared with $G_{\bf T}$ according to Definition 1. Second, if all latent variables are not intervention targets (i.e., ${\bf T}={\bf T_{O}}$), then our identifiability result is stronger than the existing results in (Jaber et al.,, 2020; Varici et al.,, 2022). In particular, Jaber et al., (2020) and Varici et al., (2022) showed that in this case, ${\bf T_{O}}$ can only be identified up to the neighbors of ${\bf N}_{{\bf T}}$ in the MAG corresponding to $G_{\bf T}$. We improve their results by adding part (a)(ii) in Definition 1, due to the recovery of $\tilde{{\bf N}}_{{\bf T}}$. See Figure 2 for a diagram comparing the recovery outputs and Example 2 below.
@@
Remark 4.
We note that LIT algorithm is capable of identifying intervention targets for each environment through extra invariance tests. In particular, for each pair of environments, we can conduct an invariance test on the samples of recovered noise in two environments.